The star of a distant solar system explodes as a supernova. At the moment of the explosion, an

resting exploration spaceship is 15 AU away from the shock wave. The shock wave of the explo-

sion travels with 25000 km/s towards the spaceship. To save the crew, the spacecra makes use

of a special booster that uniformly accelerates at 150 m/s2

in the opposite direction.

Determine if the crew manages to escape from the shock wave. (Neglect relativistic eects.)


1
Expert's answer
2021-02-15T18:15:33-0500

The time it takes for the spaceship to reach 25000 km/s:



"t=\frac{v}{a}=166667\text{ s}."

The distance the shockwave will cover during t:



"d=v_st=4.167·10^{12}\text{ m},"

or 27.85 AU.


So, only 27.85-15=12.85 AU left for the spaceship to develop the speed of 25000 km/s. Is it possible? Find the distance required for the ship to accelerate from rest to 25 000 000 m/s with acceleration of 150 m/s2:



"d=\frac{v^2}{2a}=2.08·10^{12}\text{ m},"

or 13.9 AU.

Therefore, when the shockwave will travel 27.85 AU from the star at 25 000 km/s, the speed of the spaceship will be lower than 25 000 km/s. They cannot escape.